# Get rich quick: exact solutions reveal how unbalanced initializations promote rapid feature learning

### v1

@@ 5 Piecewise Linear Networks
Unbalanced initializations in diverse domains. In our analysis, we find that upstream initializations can lead to rapid rich learning in nonlinear networks, explaining results shown in Fig. 1. Further experiments in Fig. 6 suggest that upstream initializations has an impact across various domains of deep learning: (a) Standard initializations see significant NTK evolution early in training [25]. We show the movement is linked to changes in activation patterns (Hamming distance) rather than large parameter shifts. Adjusting the initialization variance of the first and last layers can amplify or diminish this movement. (b) Filters in CNNs trained on image classification tasks often align with edge detectors [64]. We show that adjusting the learning speed of the first layer can enhance or degrade this alignment. (c) Deep learning models are believed to avoid the curse of dimensionality and learn with limited data by exploiting hierarchical structures in real-world tasks. Using the Random Hierarchy Model, introduced by Petrini et al. [65] as a framework for synthetic hierarchical tasks, we show that modifying the initialization geometry can decrease or increase the sample complexity of learning. (d) Networks trained on simple modular arithmetic tasks will suddenly generalize long after memorizing their training data [66]. This behavior, termed grokking, is thought to result from a transition from lazy to rich learning [67, 68] and believed to be important towards understanding emergent phenomena [69]. We show that decreasing the variance of the embedding in a single-layer transformer ($<6\%$ of all parameters) significantly reduces the time to grokking. Overall, our experiments suggest that upstream initializations may play a crucial role in neural network behaviors.
@@ 6 Conclusion
In this work, we derived exact solutions to a minimal model that can transition between lazy and rich learning to precisely elucidate how unbalanced layer-specific initialization variances and learning rates determine the degree of feature learning. We further extended our analysis to wide and deep linear networks and shallow piecewise linear networks. We find through theory and empirics that unbalanced initializations, which promote faster learning at earlier layers, can actually accelerate rich learning. Limitations. The primary limitation lies in the difficulty to extend our theory to deeper nonlinear networks. In contrast to linear networks, where additional symmetries simplify dynamics, nonlinear networks require consideration of the activation pattern’s impact on subsequent layers. One potential solution involves leveraging the path framework used in Saxe et al. [70]. Another limitation is our omission of discretization and stochastic effects of SGD, which disrupt the conservation laws central to our study and introduce additional simplicity biases [71, 72, 73]. Future work. Our theory encourages further investigation into unbalanced initializations to optimize efficient feature learning. In deep networks there are many ways to be unbalanced beyond upstream and downstream. Understanding how the learning speed profile across layers impacts feature learning, inductive biases, and generalization is an important direction for future work.
@@
This project originated from conversations between Daniel and Allan at the Kavli Institute for Theoretical Physics. Daniel, Allan, and Feng are primarily responsible for the single neuron analysis in Section 3. Clem and Daniel are primarily responsible for the wide and deep linear analysis in Section 4. Daniel is primarily responsible for the nonlinear analysis in Section 5. Allan, Feng, and David are primarily responsible for the empirics in Fig. 1 and Fig. 6. Daniel is primarily responsible for writing the main sections. All authors contributed to the writing of the appendix and the polishing of the manuscript.
@@ A.2.3 Upstream $\delta>0$
which can be solved analytically using $\nu(t)$. We omit the solution due to its complexity, but provide a notebook used to generate our figures encoding the solution. From the solutions for $\nu,a$, we can easily obtain dynamics for $\mu,\phi$.
@@ A.2.4 Downstream $\delta<0$
$\dot{\omega}$ $=\upsilon\|\beta\|^{2}\omega-\upsilon^{2}\omega^{3},$ $\omega(0)=w(0)^{\intercal}\beta_{*},$ (32)
which can be solved analytically using $\upsilon(t)$. Again, we omit the solution due to its complexity, but provide a notebook used to generate our figures encoding the solution. From the solutions for $\upsilon,\omega$, we can easily obtain dynamics for $\mu,\phi$.
@@ A.2.6 Recovering parameters $(a,w)$ from $(\mu,\phi)$
where $c_{1}(t)$ is the coefficient in the direction of $\beta_{*}$ and $c_{2}(t)$ is the coefficient in the direction orthogonal to $\beta_{*}$ on the two-dimensional plane defined by $w_{0}$. From the definition of $\phi$ we can easily obtain the coefficients $c_{1}=\|w\|\phi$ and $c_{2}=\sqrt{\|w\|^{2}-c_{1}^{2}}$. We always choose the positive square root for $c_{2}$, as $c_{2}(t)\geq 0$ for all $t$.
@@ A.3 Function space dynamics of $\beta$
$\dot{\beta}=a\dot{w}+\dot{a}w=-\underbrace{\left(\eta_{w}a^{2}I_{d}+\eta_{a}ww^{\intercal}\right)}_{M}\underbrace{\left(X^{\intercal}X\beta-X^{\intercal}y\right)}_{X^{\intercal}\rho}.$ (36)
@@
$M=\frac{\sqrt{\delta^{2}+4\eta_{a}\eta_{w}\|\beta\|^{2}}+\delta}{2}I_{d}+\frac{\sqrt{\delta^{2}+4\eta_{a}\eta_{w}\|\beta\|^{2}}-\delta}{2}\frac{\beta\beta^{\intercal}}{\|\beta\|^{2}}.$ (39)
Thus, given any initialization $a_{0},w_{0}$ such that $a(t)^{2}>0$ for all $t\geq 0$, we can express the dynamics of $\beta$ entirely in terms of $\beta$. This is true for all initializations with $\delta\geq 0$, except if initialized on the saddle point at the origin. It is also true for all initializations with $\delta<0$ where the sign of $a$ does not switch signs. In the next section we will show how to interpret these trajectories as time-warped mirror flows for a potential that depends on $\delta$. As a means of keeping the analysis entirely in $\beta$ space, we will make the slightly more restrictive assumption to only study trajectories given any initialization $\beta_{0}$ such that $\|\beta(t)\|>0$ for all $t\geq 0$.
@@ A.4.1 Overview of time-warped mirror flow analysis
which even for extremely simple Jacobian maps may not be true [47]. Azulay et al. [9] showed that if there exists a smooth scalar function $g(\beta):\mathbb{R}^{d}\to\mathbb{R}$ such that the ODE,
@@ A.4.2 Applying time-warped mirror flow analysis
$c_{\delta}=\begin{cases}0&\text{if }\delta\leq 0\\ \frac{\sqrt{2}|\delta|^{\frac{3}{2}}}{3}&\text{if }\delta>0\end{cases}=\max\left\{0,\mathrm{sgn}(\delta)\frac{\sqrt{2}|\delta|^{\frac{3}{2}}}{3}\right\},$ (66)
@@
$\Phi_{\delta}(\beta)=\frac{1}{3}\left(\sqrt{\delta^{2}+4\|\beta\|^{2}}-2\delta\right)\sqrt{\sqrt{\delta^{2}+4\|\beta\|^{2}}+\delta}+\max\left\{0,\mathrm{sgn}(\delta)\frac{\sqrt{2}|\delta|^{\frac{3}{2}}}{3}\right\}.$ (67)
@@ Appendix B Wide and Deep Linear Networks
In the previous section we demonstrated how the balancedness $\delta$ influences the regime of learning in a single-neuron linear network by studying the dynamics in parameter space, function space, and the implicit bias. Throughout our analysis, we identified three learning regimes – lazy, rich, and delayed rich – that correspond to different values of $\delta$. The driving cause of this distinction is the change in the geometry and topology of the conserved surface. Here we discuss how our analysis techniques can be extended to linear networks with multiple neurons, layers, and outputs. As we move towards more complex networks, the number of conserved quantities will grow, one for each hidden-neuron. As a result, the analysis in this section will get more complex, but overall the main points identified in the single-neuron setting still hold.
B.1 Two layer function space dynamics.
@@
We consider the dynamics of a two-layer linear network with $h$ hidden neurons and $c$ outputs, $f(x;\theta)=A^{\intercal}Wx$, where $W\in\mathbb{R}^{h\times d}$ and $A\in\mathbb{R}^{h\times c}$. We assume that $\min\{d,c\}\leq h\leq\max\{d,c\}$, such that this parameterization can represent all linear maps from $\mathbb{R}^{d}\to\mathbb{R}^{c}$. As in the single-neuron setting, the rescaling symmetry in this model between the first and second layer implies the $h\times h$ matrix $\Delta=A_{0}A_{0}^{\intercal}-W_{0}W_{0}^{\intercal}$ determined at initialization remains conserved throughout gradient flow [62]. The NTK matrix can be expressed as $K=\left(\mathbf{I}_{c}\otimes X\right)\left(\eta_{w}A^{\intercal}A\oplus\eta_{a}W^{\intercal}W\right)\left(\mathbf{I}_{c}\otimes X^{\intercal}\right)$, where $\otimes$ and $\oplus$ denote the Kronecker product and sum [^7] The Kronecker sum is defined for square matrices $A\in\mathbb{R}^{c\times c}$ and $B\in\mathbb{R}^{d\times d}$ as $A\oplus B=A\otimes\mathbf{I}_{d}+\mathbf{I}_{c}\otimes B$. respectively. We consider the dynamics of $\beta=W^{\intercal}A\in\mathbb{R}^{d\times c}$ in function space. The network function $\beta$ is governed by the ODE,
$\dot{\beta}=\dot{W}^{\intercal}A+W^{\intercal}\dot{A}.$ (73)
Respectively $W$ and $A$ follow the temporal dynamics given by the ODE
$\dot{W}^{\intercal}=-\eta_{w}X^{\intercal}(X\beta-Y)A^{\intercal},$ (74)
and
$\dot{A}=-\eta_{a}WX^{\intercal}(X\beta-Y).$ (75)
Replacing equations 75 and 74 in equation 73 we get
$\dot{\beta}=-\left(\eta_{w}X^{\intercal}(X\beta-Y)A^{\intercal}A+\eta_{a}W^{\intercal}WX^{\intercal}(X\beta-Y)\right)).$ (76)

### v2

@@ 5 Piecewise Linear Networks
Unbalanced initializations in practice. Our analysis shows that upstream initializations can drive rapid rich learning in nonlinear networks. Further experiments in Fig. 6 show that upstream initializations are relevant across various domains of deep learning: (a) Standard initializations see significant NTK evolution early in training [27]. We show the movement is linked to changes in activation patterns rather than large parameter shifts. Adjusting the initialization variance of the first and last layers can amplify or diminish this movement. (b) Filters in CNNs trained on image classification tasks often align with edge detectors [66]. We show that adjusting the learning speed of the first layer can enhance or degrade this alignment. (c) Deep learning models are believed to avoid the curse of dimensionality and learn with limited data by exploiting hierarchical structures in real-world tasks. Using the Random Hierarchy Model, introduced by Petrini et al. [67] as a framework for synthetic hierarchical tasks, we show that modifying the relative scale can decrease or increase the sample complexity of learning. (d) Networks trained on simple modular arithmetic tasks will suddenly generalize long after memorizing their training data [68]. This behavior, termed grokking, is thought to result from a transition from lazy to rich learning [69, 70, 71] and believed to be important towards understanding emergent phenomena [72]. We show that decreasing the variance of the embedding in a single-layer transformer ($<6\%$ of all parameters) significantly reduces the time to grokking.
@@ 6 Conclusion
In this work, we derived exact solutions to a minimal model that can transition between lazy and rich learning to precisely elucidate how unbalanced layer-specific initialization variances and learning rates determine the degree of feature learning. We further extended our analysis to wide and deep linear networks and shallow piecewise linear networks. We find through theory and empirics that unbalanced initializations, which promote faster learning at earlier layers, can actually accelerate rich learning. Limitations. The primary limitation lies in the difficulty to extend our theory to deeper nonlinear networks. In contrast to linear networks, where additional symmetries simplify dynamics, nonlinear networks require consideration of the activation pattern’s impact on subsequent layers. One potential solution involves leveraging the path framework used in Saxe et al. [73]. Another limitation is our omission of discretization and stochastic effects of SGD, which disrupt the conservation laws central to our study and introduce additional simplicity biases [74, 75, 76, 77]. Future work. Our theory encourages further investigation into unbalanced initializations to optimize efficient feature learning. Understanding how the learning speed profile across layers impacts feature learning, inductive biases, and generalization is an important direction for future work.
@@
This project originated from conversations between Daniel and Allan at the Kavli Institute for Theoretical Physics. Daniel, Allan, and Feng are primarily responsible for the single neuron analysis in Section 3. Daniel, Clem, Allan, and Feng are primarily responsible for the wide and deep linear analysis in Section 4. Daniel is primarily responsible for the nonlinear analysis in Section 5. Allan, Feng, and David are primarily responsible for the empirics in Fig. 1 and Fig. 6. Daniel is primarily responsible for writing the main sections. All authors contributed to the writing of the appendix and the polishing of the manuscript.
@@ A.2.3 Upstream $\delta>0$
which can be solved analytically using $\nu(t)$. For $a_{0}>0$, we have that
$\begin{split}a(t)=2e^{t\delta/2}\|\beta_{*}\|\sqrt{\delta}\Bigg{(}&\mathrm{sech}^{2}\left(Y(t)\right)\bigg{[}4e^{t\delta}\|\beta_{*}\|^{2}-\frac{\left(\delta^{2}+4\|\beta_{*}\|^{2}\right)\left(\|\beta_{*}\|^{2}\left(\delta-a_{0}^{2}\right)+b_{0}^{2}\right)}{b_{0}^{2}-a_{0}^{2}\|\beta_{*}\|^{2}+a_{0}b_{0}\delta}\\ &-\delta e^{\delta t}\left(\delta\cosh\left(2Y(t)\right)-\sqrt{\delta^{2}+4\|\beta_{*}\|^{2}}\sinh\left(2Y(t)\right)\right)\bigg{]}\Bigg{)}^{-1/2}\end{split}$
where $b_{0}=w_{0}^{\intercal}\beta_{*}$, and $Y(t)=\frac{1}{2}\sqrt{\delta^{2}+4\|\beta_{*}\|^{2}}t+\mathrm{atanh}\left(\frac{\frac{2b_{0}}{a_{0}}+\delta}{\sqrt{\delta^{2}+4\|\beta_{*}\|^{2}}}\right)$. From the solutions for $\nu,a$, we can easily obtain dynamics for $\mu,\phi$.
@@ A.2.4 Downstream $\delta<0$
$\dot{\omega}$ $=\omega\left(\upsilon\|\beta\|^{2}-\upsilon^{2}\omega^{2}\right),$ $\omega(0)=w(0)^{\intercal}\beta_{*},$ (33)
which can be solved analytically using $\upsilon(t)$, analogous to the solution for $a(t)$ in Section A.2.3. From the solutions for $\upsilon,\omega$, we can easily obtain dynamics for $\mu,\phi$.
@@ A.2.6 Recovering parameters $(a,w)$ from $(\mu,\phi)$
where $c_{1}(t)$ is the coefficient in the direction of $\beta_{*}$ and $c_{2}(t)$ is the coefficient in the direction orthogonal to $\beta_{*}$ on the two-dimensional plane defined by $w_{0}$. From the definition of $\phi$ we can easily obtain the coefficients $c_{1}=\|w\|\phi$ and $c_{2}=\sqrt{\|w\|^{2}-c_{1}^{2}}$. We always choose the positive square root for $c_{2}$, as $c_{2}(t)\geq 0$ for all $t$. See Section D.2 for experimental details of how we ran our simulations and a notebook generating these exact solutions.
@@ A.3 Function space dynamics of $\beta$
$\dot{\beta}=a\dot{w}+\dot{a}w=-\underbrace{\left(\eta_{w}a^{2}\mathbf{I}_{d}+\eta_{a}ww^{\intercal}\right)}_{M}\underbrace{\left(X^{\intercal}X\beta-X^{\intercal}y\right)}_{X^{\intercal}\rho}.$ (37)
@@
$M=\frac{\sqrt{\delta^{2}+4\eta_{a}\eta_{w}\|\beta\|^{2}}+\delta}{2}\mathbf{I}_{d}+\frac{\sqrt{\delta^{2}+4\eta_{a}\eta_{w}\|\beta\|^{2}}-\delta}{2}\frac{\beta\beta^{\intercal}}{\|\beta\|^{2}}.$ (40)
Thus, given any initialization $a_{0},w_{0}$ such that $a(t)^{2}\neq 0$ for all $t\geq 0$, we can express the dynamics of $\beta$ entirely in terms of $\beta$. This is true for all initializations with $\delta\geq 0$, except if initialized on the saddle point at the origin. It is also true for all initializations with $\delta<0$ where the sign of $a$ does not switch signs. In the next section we will show how to interpret these trajectories as time-warped mirror flows for a potential that depends on $\delta$. As a means of keeping the analysis entirely in $\beta$ space, we will make the slightly more restrictive assumption to only study trajectories given any initialization $\beta_{0}$ such that $\|\beta(t)\|>0$ for all $t\geq 0$.
@@
A.3.1 Kernel dynamics
The dynamics of the NTK matrix $K=XMX^{\intercal}$ is determined by $\dot{M}$. From Eq. 4, which is derived in this section, we can write $\dot{M}=\frac{2\|\beta\|}{\kappa}(\mathbf{I}_{d}+\hat{\beta}\hat{\beta}^{\intercal})\partial_{t}\|\beta\|+\frac{\kappa-\delta}{2}\partial_{t}(\hat{\beta}\hat{\beta}^{\intercal})$ where $\hat{\beta}=\tfrac{\beta}{\|\beta\|}$. From this expression we see that the change in $M$ is driven by two terms, one that depends on the change in the magnitude of $\beta$ and another that depends on the change in the direction of $\beta$. As done in the main text, we consider $\delta\gg 0$, $\delta\ll 0$, and $\delta=0$ to identify different regimes of learning. For $\delta\gg 0$, the coefficients in front of both terms vanish, and thus, irrespective of the trajectory taken from $\beta(0)$ to $\beta_{*}$, the change in the NTK is vanishing, indicative of a lazy regime. For $\delta\ll 0$, the coefficient for the first term vanishes, while the coefficient on the second term diverges. Here, the change in the NTK is driven solely by the change in the direction of $\beta$. This is why for large negative delta we observe a delayed rich regime, where the eventual alignment of $\beta$ to $\beta_{*}$ leads to a dramatic change in the kernel. When $\delta=0$, the coefficients for both terms are roughly of the same order, and thus changes in both the magnitude and direction of $\beta$ contribute to a change in the kernel, indicative of a rich regime.
@@ A.4.1 Overview of time-warped mirror flow analysis
which even for extremely simple Jacobian maps may not be true [47]. Azulay et al. [9] showed that if there exists a smooth positive function $g(\beta):\mathbb{R}^{d}\to(0,\infty)$ such that the ODE,
@@ A.4.2 Applying time-warped mirror flow analysis
$c_{\delta}=\begin{cases}0&\text{if }\delta\leq 0\\ \frac{\sqrt{2}|\delta|^{\frac{3}{2}}}{3}&\text{if }\delta>0\end{cases}=\max\left(0,\mathrm{sgn}(\delta)\frac{\sqrt{2}|\delta|^{\frac{3}{2}}}{3}\right),$ (67)
@@
$\Phi_{\delta}(\beta)=\frac{1}{3}\left(\sqrt{\delta^{2}+4\|\beta\|^{2}}-2\delta\right)\sqrt{\sqrt{\delta^{2}+4\|\beta\|^{2}}+\delta}+\max\left(0,\mathrm{sgn}(\delta)\frac{\sqrt{2}|\delta|^{\frac{3}{2}}}{3}\right).$ (68)
@@ Appendix B Wide and Deep Linear Networks
Here we discuss how our analysis techniques, developed in the previous section for a single-neuron linear network, can be extended to linear networks with multiple neurons, outputs, and layers.
B.1 Wide linear networks
@@
We consider the dynamics of a two-layer linear network with $h$ hidden neurons and $c$ outputs, $f(x;\theta)=A^{\intercal}Wx$, where $W\in\mathbb{R}^{h\times d}$ and $A\in\mathbb{R}^{h\times c}$. We assume that $h\geq\min(d,c)$, such that this parameterization can represent all linear maps from $\mathbb{R}^{d}\to\mathbb{R}^{c}$. As in the single-neuron setting, the rescaling symmetry in this model between the first and second layer implies the $h\times h$ matrix $\Delta=A_{0}A_{0}^{\intercal}-W_{0}W_{0}^{\intercal}$ determined at initialization remains conserved throughout gradient flow [62]. This can be easily shown from the temporal dynamics of $A$ and $W$,
$\dot{A}$ $=-\eta_{a}WX^{\intercal}(X\beta-Y),$ (74)
$\dot{W}^{\intercal}$ $=-\eta_{w}X^{\intercal}(X\beta-Y)A^{\intercal}.$ (75)
Extending derivations in [30], the NTK matrix can be expressed as
$K=\left(\mathbf{I}_{c}\otimes X\right)\left(\eta_{w}A^{\intercal}A\oplus\eta_{a}W^{\intercal}W\right)\left(\mathbf{I}_{c}\otimes X^{\intercal}\right),$ (76)
where $\otimes$ and $\oplus$ denote the Kronecker product and sum respectively. The Kronecker sum is defined for square matrices $C\in\mathbb{R}^{c\times c}$ and $D\in\mathbb{R}^{d\times d}$ as $C\oplus D=C\otimes\mathbf{I}_{d}+\mathbf{I}_{c}\otimes D$.
B.1.1 Parameter space dynamics
Inspired by our analysis of the single-neuron setting, we introduce two coordinate transformations to study the parameter space dynamics of a wide two-layer linear network. In both analyses we assume whitened input $X^{\intercal}X=\mathbf{I}_{d}$ and let $\eta_{a}=\eta_{w}=1$. However, we will find that the analysis of the dynamics in function space, for general unwhitened data, is more tractable.
Parameter dynamics when $c=1$. Drawing insights from our analysis of the single-neuron scenario ($h=c=1$), we might consider a combination of hyperbolic and spherical coordinate transformations to study the parameter space dynamics of a wide two-layer linear network. We consider the following two quantities for each hidden neuron $k\in[h]$:
$\mu_{k}=a_{k}\|w_{k}\|,\qquad\phi_{k}=\frac{w_{k}^{\intercal}\beta_{*}}{\|w_{k}\|\|\beta_{*}\|}.$ (77)
We will also consider a new matrix quantity $Q\in\mathbb{R}^{h\times h}$ with elements $Q_{kk^{\prime}}=\frac{w_{k}^{\intercal}w_{k^{\prime}}}{\|w_{k}\|\|w_{k^{\prime}}\|}$. The resulting dynamics for $\mu$ and $\phi$ can be entirely written in terms $\mu,\phi,\Delta$:
$\dot{\mu}$ $=\sqrt{\mathrm{Diag}(\Delta)^{2}+4\mathrm{Diag}(\mu)^{2}}\left(\phi-Q\mu\right),$ (78)
$\dot{\phi}$ $=M\mathrm{Diag}(\mu)\left((\|\beta_{*}\|^{2}-\phi^{\intercal}\mu)I_{h}+\mathrm{Diag}(\phi)Q\mu-\phi^{2}\right),$ (79)
where $M=2\left(\sqrt{\mathrm{Diag}(\Delta)^{2}+4\mathrm{Diag}(\mu)^{2}}-\mathrm{Diag}(\Delta)\right)^{-1}$. Using the conserved structure of $\Delta$ we can express $Q$ as a function of $\mu$ and $M$,
$Q=M\mu\mu^{\intercal}M-M^{1/2}\Delta M^{1/2}.$ (80)
This approach yields a coupled nonlinear dynamical system with $2h$ variables. Imposing additional assumptions on the initialization, such as permutation invariance between hidden neurons, can simplify the system of differential equations. A similar approach was used by Saad and Solla [78] to derive a set of differential equations for a soft committee machine model, capturing its online learning dynamics in a teacher-student setup, which Goldt et al. [79] extended to its generalization error dynamics.
Parameter dynamics when $c=h$. In this analysis we assume an initialization such that the conserved quantities $\Delta=\delta\mathbf{I}_{h}$, an assumption we will discuss further in Section B.1.6, and that $A$ is invertible throughout training. Let $\beta_{*}=X^{\intercal}Y$, which for whitened input, is the unique minimum of the dynamics in function space. We consider the variable $\nu=A^{-1}W\beta_{*}\in\mathbb{R}^{c\times c}$. Using the identity that $\dot{A^{-1}}=-A^{-1}\dot{A}A^{-1}$ and our assumption on $\Delta$, we find that the matrix $\nu$ evolves according to the matrix Riccati ODE,
$\dot{\nu}=\beta_{*}^{\intercal}\beta_{*}-\delta\nu-\nu^{2}.$ (81)
Additionally, consider the variable $C=A^{\intercal}A$, which evolves according to the matrix Bernoulli ODE
$\dot{C}=C(\nu+\delta\mathbf{I}_{h})+(\nu+\delta\mathbf{I}_{h})^{\intercal}C-2C^{2}.$ (82)
Taken together we have found a change of variables, analogous to the one introduced in Section A.2.3 for the single-neuron setting, that evolves according to a matrix Riccati and Bernoulli equation,
$\dot{\nu}$ $=\beta_{*}^{\intercal}\beta_{*}-\delta\nu-\nu^{2},$ $\nu(0)=A_{0}^{-1}W_{0}\beta_{*},$ (83)
$\dot{C}$ $=C(\nu+\delta\mathbf{I}_{h})+(\nu+\delta\mathbf{I}_{h})^{\intercal}C-2C^{2},$ $C(0)=A_{0}^{\intercal}A_{0}.$ (84)
However, solving this system exactly as we did in the single-neuron setting is challenging. Unless we assume that $\nu$ and $\beta_{*}^{\intercal}\beta_{*}$ share the same eigenspace – allowing us to decouple the dynamics of $\nu$ into a set of scalar Riccati equations – the system cannot be easily solved. Instead, we will find that the dynamics of the product $W^{\intercal}A$ in function space is more tractable and requires fewer assumptions.
B.1.2 Function space dynamics
We consider the dynamics of $\beta=W^{\intercal}A\in\mathbb{R}^{d\times c}$ in function space, which is governed by the ODE,
$\dot{\beta}=W^{\intercal}\dot{A}+\dot{W}^{\intercal}A=-\left(\eta_{w}X^{\intercal}(X\beta-Y)A^{\intercal}A+\eta_{a}W^{\intercal}WX^{\intercal}(X\beta-Y)\right).$ (85)